What would be the genotypes of the gametes produced from the following individuals? Give the probability...

  1. What would be the genotypes of the gametes produced from the following individuals? Give the probability in fractions or percent for each gamete.
    1. DdEe
    2. AABB
    3. AaBbDd
    4. AaBbDDEERr
Answer #1

From DdEe, 4 gametes will be produced :

DE, De, dE, de.

Probability of gametes is in the ratio:

1/4: 1/4:1/4:1:4 (. This is the probability of each gamete with respect to its genotype)

From AABB

Though 4 gametes will be produced,but they will be the same i,e AB

So here the probability will be 1/1

AaBbDd will produce following gametes:

ABD (1/8)

ABd (1/8)

AbD (1/8)

Abd (1/8)

aBD (1/8)

aBd (1/8)

abD (1/8)

abd (1/8)

An individual with AaBbDDEERr will form a total of 32 gametes

Viz:

ABDER (4/32)=1/8

ABDEr (4/32) = 1/8

AbDER (4/32)= 1/8

AbDEr (4/32)=1/8

aBDER (4/32)=1/8

aBDEr (4/32)=1/8

abDER (4/32)=1/8

abDEr (4/32)=1/8

?